# RadioDiff-Loc: Diffusion Model Enhanced Scattering Congnition for NLoS Localization with Sparse Radio Map Estimation

### v1

@@ IV-B DM-Based NLoS Localization
which approximates the MAP estimate of the emitter location. Repeated sampling yields an ensemble $\{\hat{\bm{d}}^{(k)}\}$, enabling both robust estimation and uncertainty quantification. This framework transforms the intractable NLoS localization problem into a tractable, data-driven posterior inference task. Unlike prior approaches, it (i) models $p(\bm{r}\mid\bm{d},\bm{H})$ implicitly via learned generative priors; (ii) reduces sampling cost through physically grounded, vertex-targeted RSS probes; and (iii) achieves power-invariant localization via a normalized inference pipeline. The result is a scalable and physically informed method for high-accuracy emitter localization in sparse, multipath-rich environments.
@@ V-D Comparison with Traditional Localization Methods
In the experiments, we evaluate the performance of these methods under a global random sampling strategy with a sampling rate of 3.05%, where observation points are randomly selected across the entire scene. In contrast, RadioDiff-Loc is evaluated under a more challenging setting, where observations are restricted to the upper-half region of the scene with a significantly lower sampling rate. The average localization errors (in meters) of traditional methods and RadioDiff-Loc are summarized in Table II and Fig. 3. Notably, the RadioDiff-Loc result (3.035 m) corresponds to the random sampling configuration from the Vertex + Random Sampling vs. Random Sampling experiment (see Table LABEL:tab:vertex_vs_random). In that setup, the sampling rate is approximately 0.96%, and random samples constitute about 28.95% of the total sampling points.

### v2

@@ IV-B DM-Based NLoS Localization
which approximates the MAP estimate of the emitter location. Repeated sampling yields an ensemble $\{\hat{\bm{d}}^{(k)}\}$, enabling both robust estimation and uncertainty quantification. This framework transforms the intractable NLoS localization problem into a tractable, data-driven posterior inference task. Unlike prior approaches, as is shown in Fig. 2, it (i) models $p(\bm{r}\mid\bm{d},\bm{H})$ implicitly via learned generative priors; (ii) reduces sampling cost through physically grounded, vertex-targeted RSS probes; and (iii) achieves power-invariant localization via a normalized inference pipeline. The result is a scalable and physically informed method for high-accuracy emitter localization in sparse, multipath-rich environments.
@@ V-D Comparison with Traditional Localization Methods
In the experiments, we evaluate the performance of these methods under a global random sampling strategy with a sampling rate of 3.05%, where observation points are randomly selected across the entire scene. In contrast, RadioDiff-Loc is evaluated under a more challenging setting, where observations are restricted to the upper-half region of the scene with a significantly lower sampling rate. The average localization errors (in meters) of traditional methods and RadioDiff-Loc are summarized in Table II and Fig. 4. Notably, the RadioDiff-Loc result (3.035 m) corresponds to the random sampling configuration from the Vertex + Random Sampling vs. Random Sampling experiment. In that setup, the sampling rate is approximately 0.96%, and random samples constitute about 28.95% of the total sampling points.